# Advanced Weights for IXPE Polarization Analysis

### v1

@@ 4 Application to IXPE Data
To validate weights performance with real data, we compare the polarization sensitivity of standard analyses with that of our new techniques for three previously analyzed IXPE sources. These are the bright Crab pulsar (PSR B0351++21) and its resolved PWN, the “Crab twin” (PSR B0540-69, hereafter B0540) in the large Magellanic cloud and its unresolved PWN, and GRB 221009A. IXPE observations of these have been well described in earlier publications (Bucciantini et al., 2023b; Xie et al., 2024; Negro et al., 2023, respectively). We report uniform polarizations for the central point source over a broad energy range, although each of these objects also includes more complex spatial, spectral, and temporal polarization behavior. We focus on simple sensitivity here and do not attempt detailed analysis of the signals. In the conclusions we do comment briefly on the methods’ promise for extracting more physics from IXPE data. Descriptions of the analysis for each source are given in appendix D.
@@
First on the left, we treat Mom-reconstructed data with unweighted PCUBE (Eq. 11, row 1). Polarization sensitivity generally improves when Mom event weights are added (row 2) and when background is suppressed using the cuts of Di Marco et al. (2023) (row 3). In row 4, we apply the MLE aperture analysis with background rate correction (Eqs. 3 and 10) instead of PCUBE, keeping the particle cut in place but removing Mom event weights. In all cases, this provides better sensitivity than previous rows. We conclude that the MLE-based analysis using the mean modulation factor is preferred over PCUBE Mom event-weighted analyses. To test whether Mom event weights can improve the MLE analysis, we tried a “weighted MLE analysis” (weighting the sums of Eqs. 3 and 10 with $w_{i}$). The result improves on the PCUBE weighted analysis, but still possesses larger uncertainties than the unweighted MLE by a factor of 101510-15%.
@@
On the right, we summarize results from the NN-based weighting techniques. First, NN-reconstruction with the MLE method (Eqs. 3 and 10) and energy-binned modulation factor $\mu(E_{i})$ provides improved sensitivity compared to even the best Mom-bases analysis (row 4, left) and greatly improved sensitivity compared to standard PCUBE (row 1, left). Next, row 2 shows how the sensitivity is further improved by the full event weights $\mu_{i}$. Note that, unlike Mom event weights, these do provide an improvement over MLE $\mu(E_{i})$ results, and should be used. The uncertainties for the above analytical MLE methods are obtained from linear propagation of errors.
@@
Row 3 adds spectral weights, which provide some discrimination between sources. [^2] Astrophysical spectra are obtained from former studies as reported in appendix D, the solar photon background is assigned a power law with photon index $\Gamma=2.5$ which closely matches the data, and NN- and Mom-particle background spectra are measured from data as described in appendix C. This row and those below are obtained by numerically maximizing the likelihood. Row 4 shows how PSF-weighting improves isolation of the point source from polarized background. This PSF analysis also reduces systematic error from leakage effects which occur for small or non-circular apertures. Small-aperture effects are relevant for Crab, but not quantified here. Next, additional improvement is realized by the NN particle background de-weighting. The effect is very small for these point-source extractions—the improvements should be much more substantial for measurement of faint, extended sources. Finally, to illustrate the effect of additional weights, row 5 shows how phase weighting can improve polarization measurements of the pulsar sources in wide phase windows.
@@
Overall one sees that the weighting improves our sensitivity FoM by $\gtrsim 2\times$ over simple PCUBE and by $\gtrsim 30\%$ over even the best current Mom-based analysis. A few special cases deserve comment. While MLE analysis typically offers a large gain, for B0540 we have only measured over 464-6 keV to be consistent with the original publication. In this high-energy band, the modulation factor is nearly constant, so the MLE analysis provides little gain over PCUBE. If the detection band can be widened, improvements should be larger. On the other hand, NN reconstruction performs best for these high energy tracks compared to Mom, so NN uncertainties are much reduced. Also, spatial weighting provides good gain when the source/background ratio flux is small. Crab peak 2 is a good example. As noted, background suppression weights offer modest gain for these bright point sources but can be significant for extended faint sources. Our prescription for the background de-weighting appropriately adjusts to the varying circumstances: when background is weak, particle background is only lightly cleaned (e.g. Crab). When particle background is strong, then the weights have a larger effect (e.g. the GRB).
@@ 5 Discussion and Conclusions
This paper applies spatial weights to point sources (albeit with spatially varied backgrounds), but the methods may readily be applied also to extended sources as well. An analysis of a complex field might proceed as follows: an image delivered by the Chandra X-ray observatory gives a high-resolution, unpolarized X-ray image, which the astronomer separates into possibly polarized components. The image is rescaled to the IXPE band, and spatially uniform components are added representing solar photons and particles. If necessary for a good fit, these background components should be modeled as polarized. IXPE source polarizations are extracted, using spatial, particle, phase, and/or spectral weights to separate the components. Best results are achieved using NN reconstruction and event weights.
@@
This work was supported in part by by contract NNM17AA26C from the MSFC to Stanford in support of the IXPE project. The authors also thank Niccolò Di Lalla, Luca Baldini, and the rest of the IXPE Collaboration for helpful conversations.
@@ Appendix B Polarization Leakage Forward Model
To validate our spatial weights on simulated data in §3, we construct a simulation capable of modeling the relevant effects, including the telescope response functions, PSF, and polarization leakage. While we cannot use the IXPEobssim simulator because it does not model polarization leakage, we otherwise mirror its approach. We draw event energies from power-law spectra with photon indices of 1.5 for the source and 2 for the background. The source spectrum is multiplied by the IXPE effective area and quantum efficiency as given in IXPEobssim. The EVPA $\psi$ of each event is drawn from the standard Eq. 2. The modulation factor $\mu$ is also given as a function of energy in IXPEobssim.
@@
Appendix C Neural Net Event Reconstruction
@@
Appendix D Detailed Analysis Methodology
Crab
The Crab is bright enough that a single observation (ID 02001601) supplies plenty of signal. We apply the NN-reconstruction method as described in appendix C and phase the data after barycentric correction using the CRABTIME ephemeris database. After aligning the detectors to center the pulsar, our polarization is extracted from a 25′′ aperture during both the “on” and “off” phases and subtracted to reveal the pulsar polarization. For the spatially weighted methods, we model the nebula within this aperture with a flux map taken from a contemporaneous Chandra observation (observation ID 23539), reprocessed, cut to 282-8 keV, and mapped to the IXPE effective area. Our pulsar spectral model is taken from (Weisskopf et al., 2004) and the nebula spectrum is assumed to exhibit $\Gamma=1.9$, consistent with torus spectra measured in (Mori et al., 2004). Since the polarization maps reported in (Wong et al., 2023) found little PA variation within our region, assuming uniform polarization for this structure is acceptable. Spatial variation of the nebular background will produce unphysical polarization leakage signals that will not average to zero inside the aperture. While the PSF-weighted analysis automatically corrects for this leakage, it is not corrected in the aperture extraction methods.
Our measured background (nebula) PA agrees with that previously determined for the center of the Crab (Wong et al., 2023). For the pulsar, we extract separate PDs for the first (phase 0.170.240.17-0.24) and second (0.5650.650.565-0.65) peaks. Note that more detailed phase analysis show rapid PA sweeps (Bucciantini et al., 2023a; Wong et al., 2023), so phases values are sensitive to the precise boundaries. However our results in these broad peak bins are consistent with previous measurements and more importantly use identical events for each analysis, allowing useful FoM comparisons.
PSR B0540-69
Our B0540 analysis considers IXPE 4-6 keV events from observations 02001201, 02001202, and 02008801, as the original analysis finds significant polarization only in this bin (Xie et al., 2024). When we reprocess NN events, we retain the Mom energy estimates to ensure that we fit to the same events regardless of the reconstruction method. [^3] This 2 keV range is not much larger than the IXPE energy resolution, so different event reconstruction can significantly alter the included set. We use the same ephemeris, spectrum, and on-off phase bins as that analysis. Our aperture-based methods are conducted with a radius of 60′′, and the off-phase is taken as background for the on-phase, similarly to Crab. The PSF-weighted analyses use a 278′′-radius aperture, nearly to the chip edge. Using a large aperture is beneficial because the PSF wings provide some useful counts at these large separations, and spatial weighting ensures the extra background events do not harm the fit precision. Our phase-weighted analysis uses the entire available B0540 data set. We see a slight, non-significant drop in PD which would be expected if the PA sweeps over the pulsar period.
[…]
GRB 221009A
Observation 02250101 of the B.O.A.T. was first analyzed using a 26′′ aperture surrounding the point source (Negro et al., 2023). We use the same radius for aperture-based methods, and assign spectral weights using the spectrum found in that work. Lacking a clean background region, we approximate background flux using a $160-200^{\prime\prime}$-radius annulus, which will be slightly polluted by flux from the scattered rings (see Fig. 1). We do not require a clean background region for spatially weighted methods, so for these we remove the background annulus and extend the source aperture to 60′′. In principle, this source aperture extension risks including some flux from the extended rings of scattered X-rays from the initial explosion. However, at the time of the IXPE observation, the the dust responsible for ring emission within our aperture is at distances >48>48 kpc from Earth, where densities are likely quite low (see appendix A). Indeed, a fit that includes a ring component shows that its flux within our aperture is consistent with zero. We therefore ignore the rings in the final spatial weighted fit. We also fit uniform background photon and particle components. As with B0540, we see no signal of background polarization and therefore assume that they are unpolarized.

### v2

@@ 4 Application to IXPE Data
To validate weights performance with real data, we compare the polarization sensitivity of standard analyses with that of our new techniques for three previously analyzed IXPE sources. These are the bright Crab pulsar (PSR B0351++21) and its resolved PWN, the “Crab twin” (PSR B0540-69, hereafter B0540) in the large Magellanic cloud and its unresolved PWN, and GRB 221009A. IXPE observations of these have been well described in earlier publications (Bucciantini et al., 2023b; Xie et al., 2024; Negro et al., 2023, respectively). We report uniform polarizations for the central point source over a broad energy range, although each of these objects also includes more complex spatial, spectral, and temporal polarization behavior. We discuss only polarization sensitivity here and do not attempt detailed analysis of the signals. In the conclusions we do comment briefly on the methods’ promise for extracting more physics from IXPE data. Descriptions of the analysis for each source are given in appendix C.
@@
First on the left, we treat Mom-reconstructed data with unweighted PCUBE (Eq. 12, row 1). Polarization sensitivity generally improves when Mom event weights are added (row 2) and when background is suppressed using the cuts of Di Marco et al. (2023) (row 3). In row 4, we apply the MLE aperture analysis with background rate correction (Eqs. 3 and 11) instead of PCUBE, keeping the particle cut in place but removing Mom event weights. In all cases, this provides better sensitivity than previous rows. We conclude that the MLE-based analysis using the mean modulation factor is preferred over PCUBE Mom event-weighted analyses. To test whether Mom event weights can improve the MLE analysis, we tried a “weighted MLE analysis” (weighting the sums of Eqs. 3 and 11 with $w_{i}$). The result improves on the PCUBE weighted analysis, but still possesses larger uncertainties than the unweighted MLE by a factor of 101510-15%.
@@
On the right, we summarize results from the NN-based weighting techniques. First, NN-reconstruction with the MLE method (Eqs. 3 and 11) and energy-binned modulation factor $\mu_{E_{i}}$ provides improved sensitivity compared to even the best Mom-bases analysis (row 4, left) and greatly improved sensitivity compared to standard PCUBE (row 1, left). Next, row 2 shows how the sensitivity is further improved by the full event weights $\mu_{i}$. Note that, unlike Mom event weights, these do provide an improvement over MLE $\mu_{E_{i}}$ results, and should be used. The uncertainties for the above analytical MLE methods are obtained from linear propagation of errors.
@@
Row 3 adds spectral weights, which provide some discrimination between sources. [^3] Astrophysical spectra are obtained from former studies as reported in appendix C, the solar photon background is assigned a power law with photon index $\Gamma=2.5$ which closely matches the data, and NN- and Mom-particle background spectra are measured from data as described in appendix D. This row and those below are obtained by numerically maximizing the likelihood. Row 4 shows how PSF-weighting improves isolation of the point source from polarized background. This PSF analysis also reduces systematic error from leakage effects which occur for small or non-circular apertures. Small-aperture effects are relevant for Crab, but not quantified here. Next, additional improvement is realized by the NN particle background de-weighting. The effect is very small for these point-source extractions—the improvements should be much more substantial for measurement of faint, extended sources. Finally, to illustrate the effect of additional weights, row 5 shows how phase weighting can improve polarization measurements of the pulsar sources in wide phase windows.
Overall one sees that the weighting improves our sensitivity FoM by $\gtrsim 2\times$ over basic PCUBE, the most common IXPE analysis method. The best Mom-based analysis, labeled MLE + ptcl. cut, has been used in several previous IXPE studies. Our analysis yields a $\gtrsim 30\%$ improvement over that combination. NN analysis has been previously presented for simulated data, so MLE, NN data + NN evt. wts could have been previously used. However, the NN analysis has not yet been applied to real IXPE data in peer-reviewed publications, so we include it as a new weight here.
@@
A few special cases deserve comment. While MLE analysis typically offers a large gain, for B0540 we have only measured over 464-6 keV to be consistent with the original publication. In this high-energy band, the modulation factor is nearly constant, so the MLE analysis provides little gain over PCUBE. If the detection band can be widened, improvements should be larger. On the other hand, NN reconstruction performs best for these high energy tracks compared to Mom, so NN uncertainties are much reduced. Also, spatial weighting provides good gain when the source/background ratio flux is small. Crab peak 2 is a good example. As noted, background suppression weights offer modest gain for these bright point sources but can be significant for extended faint sources. Our prescription for the background de-weighting appropriately adjusts to the varying circumstances: when background is weak, particle background is only lightly cleaned (e.g. Crab). When particle background is strong, then the weights have a larger effect (e.g. the GRB).
@@ 5 Discussion and Conclusions
This paper applies spatial weights to point sources (albeit with spatially varied backgrounds), but the methods may readily be applied also to extended sources as well. An analysis of a complex field might proceed as follows: an image delivered by the Chandra X-ray observatory gives a high-resolution, unpolarized X-ray image, which the astronomer separates into possibly polarized components. The image is rescaled to the IXPE band, and spatially uniform components are added representing solar photons and particles. If necessary for a good fit, these background components should be modeled as polarized. IXPE polarizations for each source are extracted, using spatial, particle, phase, and/or spectral weights to separate the components. Best results are achieved using NN reconstruction and event weights.
@@
This work was supported in part by contract NNM17AA26C from the MSFC to Stanford in support of the IXPE project. The authors also thank Niccolò Di Lalla, Luca Baldini, and the rest of the IXPE Collaboration for helpful conversations. During the review process, the authors became aware of Ravi et al. (2025), which presents a Bayesian analysis tool (QUEEN-BEE) for polarization analysis of a time-varying source. In the special case of a bright point source, the two methods use the same likelihoods and priors and should give identical results. Both methods present likelihood-based analyses as the future of IXPE polarimetry.
@@ Appendix B Polarization Leakage Forward Model
To validate our spatial weights on simulated data in §3, we construct a simulation capable of modeling the relevant effects, including the telescope response functions, PSF, and polarization leakage. While we cannot use the IXPEobssim simulator because it does not model polarization leakage, we otherwise mirror its approach. We draw event energies from power-law spectra with photon indices of $\Gamma=1.5$ for the source and 2 for the background. The source spectrum is multiplied by the IXPE effective area and quantum efficiency as given in IXPEobssim. The EVPA $\psi_{i}$ of each event is drawn from the standard Eq. 2 using the modulation factor $\mu_{E_{i}}$ given in IXPEobssim.
@@
Appendix C Detailed Analysis Methodology
Crab
The Crab is bright enough that a single observation (ID 02006001) supplies plenty of signal. We apply the NN-reconstruction method as described in appendix D and phase the data after barycentric correction using the CRABTIME ephemeris database. After aligning the detectors to center the pulsar, our polarization is extracted from a 25′′ aperture during both the “on” and “off” phases and subtracted to reveal the pulsar polarization. For the spatially weighted methods, we model the nebula within this aperture with a flux map taken from a contemporaneous Chandra observation (observation ID 23539), reprocessed, cut to 282-8 keV, and mapped to the IXPE effective area. For energy weights, our pulsar spectral model is taken from (Weisskopf et al., 2004) and the nebula spectrum is assumed to exhibit $\Gamma=1.9$, consistent with torus spectra measured in (Mori et al., 2004). For phase weights, we measure the light curve from the region of interest and ascribe all the pulsed flux to the pulsar. Since the polarization maps reported in (Wong et al., 2023) found little PA variation within our region, assuming uniform polarization for this structure is acceptable. Spatial variation of the nebular background will produce unphysical polarization leakage signals that will not average to zero inside the aperture. The PSF-weighted analysis automatically corrects for this leakage as demonstrated with Eq. 8, but the aperture extraction methods do not.
Our measured background (nebula) PA agrees with that previously determined for the center of the Crab (Wong et al., 2023). For the pulsar, we extract separate PDs for the first (phase 0.170.240.17-0.24) and second (0.5650.650.565-0.65) peaks. Note that more detailed phase analysis shows rapid PA sweeps (Bucciantini et al., 2023a; Wong et al., 2023), so values are sensitive to the precise boundaries. However our results in these broad bins are consistent with previous measurements and more importantly use identical events for each analysis, allowing useful FoM comparisons.
PSR B0540-69
Our B0540 analysis considers IXPE 4-6 keV events from observations 02001201, 02001202, and 02008801, as the original analysis finds significant polarization only in this bin (Xie et al., 2024). When we reprocess NN events, we retain the Mom energy estimates to ensure that we fit to the same events regardless of the reconstruction method. [^4] This 2 keV range is not much larger than the IXPE energy resolution, so different event reconstruction can significantly alter the included event set. We use the same ephemeris, spectrum, and on-off phase bins as that analysis. Our aperture-based methods are conducted with a radius of 60′′, and the off-phase is taken as background for the on-phase, similarly to Crab. The PSF-weighted analyses use a 278′′-radius aperture, nearly to the chip edge. Using a large aperture is beneficial because the PSF wings provide some useful counts at these large separations, and spatial weighting ensures the extra background events do not harm the fit precision. For phase weight modeling, we measure the observation’s light curve and ascribe all the pulsed flux to the pulsar, plus a DC component with amplitude 8% of the pulse height (Xie et al., 2024). We see a slight, non-significant drop in PD which would be expected if the PA sweeps over the pulsar period.
[…]
GRB 221009A
Observation 02250101 of the B.O.A.T. was first analyzed using a 26′′ aperture surrounding the point source (Negro et al., 2023). We use the same radius for aperture-based methods, and assign spectral weights using the spectrum found in that work. Lacking a clean background region, we approximate background flux using a $160^{\prime\prime}-200^{\prime\prime}$-radius annulus, which will be slightly polluted by flux from the scattered rings (see Fig. 1). We do not require a clean background region for spatially weighted methods, so for these we remove the background annulus and extend the source aperture to 60′′. In principle, this source aperture extension risks including some flux from the extended rings of scattered X-rays from the initial explosion. However, at the time of the IXPE observation, the dust responsible for ring emission within our aperture is at distances >48>48 kpc from Earth, where densities are likely quite low (see appendix A). Indeed, a fit that includes a ring component shows that its flux within our aperture is consistent with zero. We therefore ignore the rings in the final spatial weighted fit. We also fit uniform background photon and particle components. As with B0540, we see no signal of background polarization and therefore assume that they are unpolarized.
Appendix D Neural Net Event Reconstruction